OpenStudy (anonymous):

By joining dots together, how many squares can you draw on a grid measuring 4 dots by 4 dots? what about a 5x5? Can you generalise for any size grid?

OpenStudy (anonymous):

This is again just a challenge problem for those how wish to think

OpenStudy (anonymous):

I am not sure about the generalisation for this :-)

OpenStudy (anonymous):

\[s^2+(s-1)^2+(s-2)^2...(s-s)^2\]

OpenStudy (anonymous):

4x4 ----13 squares

OpenStudy (anonymous):

4*4= 20 squares

OpenStudy (anonymous):

Oh oops sorry

OpenStudy (anonymous):

5x5 ------ 24

OpenStudy (anonymous):

it can be like this: . . . . or . . . .

OpenStudy (anonymous):

I guess you miss all the sideways ones harkirat

OpenStudy (anonymous):

there is 1 4*4, 6 3*3, 13 2*2

OpenStudy (anonymous):

So the formula for the normal oriented ones would be \[(s-1)^2+(s-2)^2...(s-s)^2\] Where s is the number of dots per side

OpenStudy (anonymous):

9+3+2+1=15 not right..

OpenStudy (anonymous):

wow 9+4+1+0=13

OpenStudy (anonymous):

Lol hold you horses

OpenStudy (anonymous):

yeah that is right, I see, than you need to add the sideways

OpenStudy (anonymous):

I can't see how to get 20 squares though

OpenStudy (anonymous):

yes i missed the sideways ones !!!

OpenStudy (anonymous):

for the 4*4 there is 2 3*3 and 4 2*2 sideways

OpenStudy (anonymous):

how much is there for the 5*5?

OpenStudy (anonymous):

well there are 20 :)

OpenStudy (anonymous):

Oh wait I got 20 sorry

OpenStudy (anonymous):

for the 5*5 there are 9 2*2 sideways, 4 3*3 and 2 4*4

OpenStudy (anonymous):

so it will be (s-2)^2+(s-3)^2 +2 sideways

OpenStudy (anonymous):

you can rotate it

OpenStudy (anonymous):

ok Andras, here's one for u A man buys 100 metal balls from a store. They r all the same shape and size and packed 10/bag. BUT by mistake one bag contains balls which are all less by the same amount of weight as compared to the weight desired and available in other bags. Shipment has to move out in 2 minutes when this discrepancy is noticed-- from the total weight of the shipment. However, an employee wins a reward from the boss by finding out the wrong bag by just one weighing (and it is not by fluke). Can u explain how ?????

OpenStudy (anonymous):

I posted a new one, now I go and make myself some food

OpenStudy (anonymous):

alright good question andras

OpenStudy (anonymous):

thanks :-)

OpenStudy (anonymous):

This is not the answer but it holds true through s=5 \[(s-1)^2+(s-2)^2...+(s-s)^2\] +\[(s-2)^2+(s-3)^2...+(s-(s+1))^2+(s-2)\]